OpenStudy (anonymous):

solve x1= -1/4 and x2=3.

OpenStudy (anonymous):

what you need to solve? it's already solved :D

OpenStudy (anonymous):

it is?? oh i think i have to turn it into a quadratic equation lmao

OpenStudy (anonymous):

If you have the two roots x1= -1/4 and x2=3m you can "back track" to find the solution. (x+1/4)*(x-3)=0 gives you x1=-1/4, x2=3. If you enter the values for x1 in the left most part of the equation you get (-1/4+1/4) which is zero. That turns the whole equation to zero. If you expand the expression (x+1/4)*(x-3) you get \[x ^{2}-(11/4)x-3/4\]. And there you have your quadratic equation. Hope this helps! //Ali